A stock price is currently $50. It is known that at the end of 6 months it will be either $45 or $55. The risk-free interest rate is 10% per annum with continuous compounding. What is the value of a 6-month European put option with a strike price of $50?
Answer #1
Upmove (U)= High price/current price=55/50=1.1
Down move (D)= Low price/current price=45/50=0.9
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.1*0.5)-0.9)/(1.1-0.9)=0.75636
Put option payoff at high price (payoff H)
=Max(Strike price-High price,0)
=Max(50-55,0)
=Max(-5,0)
=0
Put option payoff at low price (Payoff L)
=Max(Strike price-low price,0)
=Max(50-45,0)
=Max(5,0)
=5
Price of Put option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.1*0.5)*(0.756355*0+(1-0.756355)*5)
=1.16